OpenStudy (anonymous):

Sorry, one small doubt: \[\frac{2}{T}\int\limits_{0}^{\frac{T}{2}} \sin( \omega t) \sin (k \omega t) \cdot dt = \frac{1}{T} \int\limits_{0}^{\frac{T}{2}}\cos( \omega t(n-1)) - \cos( \omega t(n+1)) \cdot dt\] How this integral leads to : = \(\begin{cases} \frac{1}{2} , && n = 1 && \\0, && n \ge 2, && n \in \mathbb{N} \end{cases}\)

OpenStudy (anonymous):

\[\omega = \frac{2 \pi}{T}\]

OpenStudy (anonymous):

Let us put \(n = 1\) directly..

OpenStudy (anonymous):

First term becomes \(1\)

OpenStudy (anonymous):

Second term becomes: \[\cos(2 \omega t) = \cos(\frac{4 \pi t}{T})\]

OpenStudy (anonymous):

Where I am misinterpreting it?? :(

OpenStudy (gorv):

integral cos(w(n-1)*t) = sin(w(n-1)*t)/(w(n-1)) limit

OpenStudy (anonymous):

If you will integrate it, then you will get all Integral to be \(0\)..

OpenStudy (gorv):

yeah

OpenStudy (gorv):

for n>=2 it is 0

OpenStudy (gorv):

u have prob at n=1??

OpenStudy (anonymous):

Not for \(n=1\) ??

OpenStudy (gorv):

integrate it

OpenStudy (anonymous):

Wait.. :)

OpenStudy (gorv):

sin(4pi*t/T)/(4pi/T)

OpenStudy (anonymous):

\[\frac{1}{T} \int\limits\limits_{0}^{\frac{T}{2}}\cos( \omega t(n-1)) - \cos( \omega t(n+1)) \cdot dt = \frac{1}{T} [\frac{\sin(\omega t(n-1))}{ \omega (n-1)} - \frac{\sin(\omega t(n+1))}{ \omega (n+1)}]_{0}^{T/2}\]

OpenStudy (anonymous):

\[ \frac{1}{T} [\frac{\sin(\omega t(n-1))}{ \omega (n-1)} - \frac{\sin(\omega t(n+1))}{ \omega (n+1)}]_{0}^{T/2}\]

OpenStudy (anonymous):

Is it?

OpenStudy (xapproachesinfinity):

eh question why did you go from k to n? in your first integral

OpenStudy (gorv):

yeah ....go further

OpenStudy (gorv):

put limits

OpenStudy (anonymous):

Sorry, that is not \(k\) but \(n\).. you can take \(k=n\)..

OpenStudy (gorv):

for k=1 put in start k=1

OpenStudy (gorv):

sinwt*sinwt

OpenStudy (gorv):

sin^2wt =1-cos2wt

OpenStudy (gorv):

integral t-sin2wt/2w

OpenStudy (xapproachesinfinity):

eh okay!

OpenStudy (gorv):

\[\frac{ T }{ 2 }-\frac{ \sin \frac{ 2*\pi }{ T} \frac{ T }{ 2 }}{ \frac{ 2*\pi }{ T } }\]

OpenStudy (gorv):

sin term will be zero

OpenStudy (gorv):

so we left with T/2 and 2?t is outside integral

OpenStudy (gorv):

\[\frac{ 2 }{ T }*\frac{ T }{ 2 }=1\]

OpenStudy (gorv):

oh sorry paaji

OpenStudy (anonymous):

\[\frac{1}{T} [\frac{\sin(\omega t(n-1))}{ \omega (n-1)} - \frac{\sin(\omega t(n+1))}{ \omega (n+1)}]_{0}^{T/2} = \frac{1}{ T \omega}[\frac{\sin( \pi (n-1))}{n-1} - \frac{\sin( \pi (n+1))}{n+1}]\]

OpenStudy (gorv):

cos2x=1-2sin^2x 2sin^2x=1-cos2x

OpenStudy (gorv):

so 2 we took inside in the start ....i missed that

OpenStudy (gorv):

so ans = 1/2

OpenStudy (anonymous):

@gorv

OpenStudy (anonymous):

I want to get this result from Integration result not from looking at the integral and using your brain.. :P

OpenStudy (gorv):

well u yourself said let n=1 put it first

OpenStudy (gorv):

take the intergral further

OpenStudy (gorv):

now we need to place the value for n agree??

OpenStudy (gorv):

so at n=1 n-1 =0

OpenStudy (gorv):

n-1 is in denominator of one term

OpenStudy (gorv):

so u can not put dierctly n=1 u need to apply limit

OpenStudy (gorv):

term with denominator ( n+1) at n=1 is =0

OpenStudy (gorv):

\[\lim_{n \rightarrow 1}\frac{ 1 }{ T*w } \frac{ \sin (\pi*(n-1) }{ n-1 }\]

OpenStudy (gorv):

w=2pi/T wT=2*pi

OpenStudy (gorv):

\[\frac{ 1 }{ 2 } \lim_{n \rightarrow 1}\frac{ \sin(\pi*(n-1) }{ \pi*(n-1) }\]

OpenStudy (gorv):

=1/2 hence we also proved it by integral

OpenStudy (gorv):

\[\lim_{\theta \rightarrow 0}\frac{ \sin \theta }{ \theta }=1\]

OpenStudy (gorv):

@waterineyes

OpenStudy (anonymous):

Oh, my mind was not thinking of this.. :) Thanks paaji.. :)

OpenStudy (gorv):

welcome paaji :)